A ball is thrown horizontally from a height of 19.1 m and hits the ground with a speed that is 5.00 times its initial speed. What is the speed in the vertical direction just before the ball hits the ground?
Answer #1

let, initial velocity = v

using energy conservation,

KEi + PEi = KEf + PEf

(1/2)*m*v^2 + m*g*h = (1/2)*m*(5v)^2 + 0

v = g*h / sqrt(12)

v = 9.8*19.1 / sqrt(12)

v = 3.95 m/s

let us assume that the final velocity has an angle \Theta with horizontal.

horizontal component of final velocity will be equal to initial velocity because no external force is there

5v*cos\Theta = v

\Theta = 78.46 deg

so speed in the vertical direction is

vy = 5v*sin\Theta

vy = 5*3.95*sin(78.46)

vy = 19.35 m/s

answer